OpenStudy (anonymous):

Captain Two-Pegs is going to load his pirate ship with cannons. He has room for 10 cannons. He has allotted for 111 pounds of cannon on his ship. He can choose from 9-pound cannons (the variable n) or 12-pound cannons (the variable t). Which equations would be a constraint to Captain Two-Pegs maximizing his cannon load? n - t = 10 n = 10 + t 9n = 111 - 12t 9n = 111 + 12t

OpenStudy (anonymous):

naturally it would be 9n + 12t = 111 But manipulating this equation gives you one of the answers.

OpenStudy (anonymous):

@wio thank you! that makes so much since! can u help me with more? How can x2 = 2x + 4 be set up as a system of equations? y = x2 - 2x + 4 y = x2 + 2x + 4 y = x2 y = x2 + 2x + 4 y = x2 + 2x y = x2 + 4 y = x2 y = 2x + 4